The ionisation energy of hydrogen atom is 13.6 eV, the ionisation energy of helium atom would be


        (a)    13.6eV


(b)    27.2eV


        (c)    6.8eV


        (d)    54.4eV


Grade:11

3 Answers

Amit Saxena
10 years ago

(d)

E propoertional Z2 and Z for helium = 2

:. (E)He = 4 × 13.6 = 54.4 eV

Kamaljit
6 years ago
First Ionisation potential (energy) of He is ~2372 kJ/mol ~ 24 eV (the highest among all the elements) while the second IP is (calculated from Bohr theory) ~5250 kJ/mol ~ 54 eV.
Abhay
6 years ago
AsE=(13.6ev Z^2)÷n^2Atomic weigh of helium is 4E=13.6×4=54.4 Thanks for seen🙌🙋................